# ESTIMATION OF FIRST RETURNING SPEED IN NULL RECURRENT CONTINUOUS-TIME MARKOV CHAINS

### v1

@@ 2 Preliminaries on continuous-time Markov chains
The transition rates of a CTMC are often described using a density matrix denoted by $Q=(q_{xy})$. For a given transition probability matrix $(P(t,x,y))$, Q-matrix $Q=(q_{xy})$ is deduced by
@@
where $0\leq q_{xy}<\infty$ for $x\neq y$ and $0<q_{x}:=-q_{xx}\leq\infty,\,\sum_{y\in I}q_{xy}\geq 0.$ Throughout this paper, we consistently assume that the Q-matrix $Q=(q_{xy})$ is regular. A regular Q-matrix is characterized by being totally stable, conservative and uniquely determining a CTMC. This regularity condition implies that for $x\in I$, $q_{x}<\infty$, $\sum_{y\in I}q_{xy}=0.$ Given a Q-matrix $Q=(q_{xy})$, the uniqueness of the CTMC is a hard problem. After the first flying time, the properties of the process after reaching the boundary are not determined by the Q-matrix. Under different boundary conditions, discussions on the constructions and the dynamical properties of the process can refer to Wang1992Birth; norris1998markov; chen2004markov. Here, we simply provide a practical sufficient condition for uniqueness. According to Section 2.3 in Wang1992Birth, if
$\sup_{x\in I}q_{x}<\infty,$ (2.1)
then the CTMC is unique.
@@
Let $T_{y}=\inf\{t>0:X_{t}=y\}$ be the first hitting time of state $y$. Notice that, if the process $X_{t}$ starts from $y$, then the first hitting time $T_{y}=0$. So we need to define the first returning time. Let the holding time be $\eta:=\inf\{t>0:X_{t}\neq X_{0}\}$, then the first returning time of state $x$ is denoted by $\tau_{x}=\inf\{t>\eta:X_{t}=X_{0}=x\}$. According to Theorem II.15.4 in chung1967markov, both random variables, $T_{y}$ and $\tau_{x}$, have continuous probability density functions.
@@
For a irreducible CTMC $\{X_{t},t\geq 0\}$ with state space $I$, the following statements are equivalent.
@@
    $X_{t}$ is recurrent and the invariant measure $m(\{x\})$ satisfies $m(I)=\infty$.
@@
Condition (C): $w(t)$ is a non-negative function on $0,+\infty)$ and satisfies
$w^{\prime}(t)\geq 0,\,\,\lim_{t\rightarrow\infty}w^{\prime}(t)=0,\,\,w^{\prime\prime}(t)<0\mbox{ on }[t_{0},+\infty)\mbox{ for some }t_{0}>0.$ (C)
For example, $w(t)=1\vee\log t$, $w(t)=t^{a}(0<a<1)$, $\cdots$ Intuitively speaking, a function that satisfies Condition (C) is an increasing function whose growth rate is slower than that of a linear function. Inspired by Zhao2006null, the first returning time can be characterized more precisely by considering $E^{x}(w(\tau_{x}))$ for different weight functions $w(\cdot)$ satisfying Condition (C).
@@ 3 The first returning speed of CTMCs
For $\forall a>0,\forall T>0,$ if a non-negative function $f(t)$ is integrable on $0,+\infty)$ and satisfies that
$\int_{a}^{a+T}f(t)dt\leq\int_{0}^{T}f(t)dt,$
then $\forall T>0$, we have
@@
By Lemma (3.1), we can give the following propositions.
@@
    Thus, by Lemma (3.1), the proof is completed.
@@
    By Lemma (3.1), we obtain the second proposition.
@@
We write $f(t)\sim g(t)$ provided that $\exists C_{2}>C_{1}>0$, such that for $t$ large $C_{1}\leq\frac{f(t)}{g(t)}\leq C_{2},$ where $C_{1},C_{2}$ are independent of $t$. For a non-negative continuous function $f(\cdot)$ on $0,+\infty)$, we write $\int^{\infty}f(t)dt<\infty,$ if $\exists k>0$ such that $\int^{\infty}_{k}f(t)dt<\infty.$ If there exists a constant $C\neq 0$ such that $\lim\limits_{t\to+\infty}\frac{f(t)}{g(t)}=C$, then the convergence properties of $\int^{\infty}f(t)dt$ and $\int^{\infty}g(t)dt$ are the same. For such two integrals, we write $\int^{\infty}f(t)dt\sim\int^{\infty}g(t)dt.$
The most important result in this section will be presented. We will utilize the propsitions above to give an estimate of $E^{x}(w(\tau_{x}))$, where $w(\cdot)$ is a function satisfying Condition (C).
@@
By Theorem 3.3, for $k>0$,
$\int_{k}^{\infty}\frac{t^{a-1}}{G_{1/t}(x,x)}dt$ $=\int_{k}^{\infty}t^{a-1}\sqrt{\dfrac{1}{t}(\dfrac{1}{t}+4)}dt$
$\sim\int^{\infty}t^{a-1}\dfrac{2}{\sqrt{t}}dt.$
Hence, $E^{x}(\tau_{x}^{a})<\infty$ if and only if $a\in(0,1/2)$.
@@
Consider a CTMC with the state space $I=\mathbf{Z}^{2}$. Each state $(m,n)$ transitions to the four adjacent lattice points $(n\pm 1,m)$ and $(n,\pm m)$ at a rate of $1$. it is easy to check the process is null recurrent. The embedded chain of this process is two-dimensional simple random walk.
@@
If $w(t)=t^{a}$ and $0<a<1$, then the integral in (3.5) is always divergent. Assume that $w(t)=\log^{a}t$ where $a>0$. By Theorem 3.3, consider that
@@
$=a(1-a)\pi\int^{\infty}\dfrac{(\log t)^{a-3}}{t}dt+a\pi\int^{\infty}\dfrac{(\log t)^{a-2}}{t}dt.$
Hence, $E^{x}(\log^{a}(\tau_{x}))<\infty\Leftrightarrow\int^{\infty}\dfrac{1}{t(\log t)^{2-a}}dt<\infty\Leftrightarrow a\in(0,1)$.
@@ 4 Reversible CTMCs and electric networks
Given a network $(I,E,C)$ and two boundaries $a,b\in I$, a voltage $v(x)$ is a function $v:I\rightarrow\mathbf{R}$ that is harmonic at any $x\in I-\{a,b\}$. We say $v(x)$ as the unit voltage between $a$ and $b$, that is, $v$ is the voltage, and $v(a)=1,v(b)=0.$
@@
$P_{\Delta}(x,y)=\begin{cases}\dfrac{q_{x}}{q_{x}+\lambda}\dfrac{c_{xy}}{c_{x}},&(x,y\neq\Delta)\\ \dfrac{\lambda}{q_{x}+\lambda}.&(x\neq\Delta,y=\Delta)\end{cases}$
@@ 5 The first returning speed of birth-death process
$m_{0}=1,\,m_{n}=\dfrac{\beta_{0}\beta_{1}\beta_{2}\cdots\beta_{n-1}}{\alpha_{1}\alpha_{2}\cdots\alpha_{n-1}\alpha_{n}}(n\geq 1),$ (5.1)
@@
$s(0)=0,\,s(1)=\beta_{0}^{-1},\,s(n)=\beta_{0}^{-1}+\sum_{i=1}^{n-1}\dfrac{\alpha_{1}\alpha_{2}\cdots\alpha_{i}}{\beta_{0}\beta_{1}\cdots\beta_{i}}(n\geq 2).$ (5.2)
@@
$c_{n,n+1}=m_{n}\beta_{n}=\begin{cases}\beta_{0}&(n=0),\\ \dfrac{\beta_{0}\beta_{1}\beta_{2}\cdots\beta_{n}}{\alpha_{1}\alpha_{2}\cdots\alpha_{n-1}\alpha_{n}}&(n\geq 1),\end{cases}$ (5.3)
@@
Since what we need to estimate is the situation when $\lambda$ is small, we can always assume $a_{0}\geq b_{0}$. For the electric network corresponding to a recurrent process, since the conductance $A_{n}$ should tend to 0, we have $\lim_{n\to\infty}A_{n}<\lim_{n\to\infty}B_{n}$. By Theorem 4.8, it can be seen that the closer $A_{n-1}$ is to $B_{n}$, the better the estimation of $C_{0\leftrightarrow\Delta}(\lambda)$ will be. Using this method, we present the estimation of some important models.
@@
Assume that $\{X_{t},t\geq 0\}$ is a null-recurrent birth-death process on $\mathbf{N}$. Let $F(n)=M(n)s(n)$ and
@@
If $w(t)$ satisfies Condition (C), then $E(w(\tau))<\infty$ if and only if
$\int^{\infty}t|w^{\prime\prime}(t)|\min\left(\dfrac{1}{s(F^{-1}(t)-1)},\dfrac{M(F^{-1}(t))}{t}\right)dt<\infty.$
@@
Take $N=\inf\{n\geq 0:A_{n-1}\leq B_{n}\}$. We have
@@
$\dfrac{A_{N-1}}{2}\leq[A_{N-1},B_{N}]\leq C_{0\leftrightarrow\Delta}(\lambda)\leq A_{N-1}+B_{N}\leq 2B_{N}$
and
@@
$C_{0\leftrightarrow\Delta}\left(\dfrac{1}{t}\right)\sim\min\left(B_{N},A_{N-2}\right)=\min\left(\dfrac{M(F^{-1}(t))}{t},\dfrac{1}{s(F^{-1}(t)-1)}\right).$
Therefore, by Theorem 4.4, $E(w(\tau))<\infty$ if and only if
$\int^{\infty}t|w^{\prime\prime}(t)|\min\left(\dfrac{1}{s(F^{-1}(t)-1)},\dfrac{M(F^{-1}(t))}{t}\right)dt<\infty.$
For some special birth-death processes, Theorem 5.1 has the following corollaries.
Corollary 5.2.
Assume that $\{X_{t},t\geq 0\}$ is a null-recurrent birth-death process on $\mathbf{N}$ and $w(t)$ satisfies Condition (C). If
$\limsup_{n\rightarrow\infty}\dfrac{M(n)}{M(n-1)}<\infty,$
then $E(w(\tau))<\infty$ if and only if
$\int^{\infty}|w^{\prime\prime}(t)|M(F^{-1}(t))dt<\infty.$
Proof.
In the proof of Theorem 5.1, we may observe that
$\dfrac{\lambda M(N-1)}{2}=\dfrac{B_{N-1}}{2}\leq C_{0\leftrightarrow\Delta}(\lambda)\leq 2B_{N}=2\lambda M(N).$
@@
then we have
$C_{0\leftrightarrow\Delta}\left(\dfrac{1}{t}\right)\sim B_{N}=\dfrac{M(F^{-1}(t))}{t}.$
By Theorem 4.4, the corollary holds. ∎
Similarly, we can obtain the following corollary.
Corollary 5.3.
Assume that $\{X_{t},t\geq 0\}$ is a null-recurrent birth-death process on $\mathbf{N}$ and $w(t)$ satisfies Condition (C). If
$\limsup_{n\rightarrow\infty}\dfrac{s(n)}{s(n-1)}<\infty,$
then $E(w(\tau))<\infty$ if and only if
$\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{s(F^{-1}(t))}dt<\infty.$
Remark 5.4.
A birth-death process is said to be in natural scale if its scale function satisfies $s(n)=n$. For a birth-death process in natural scale, some properties of local martingales will also be possessed. We can use Corollary 5.3 to estimate the first returning time of this type of process.
Here, we present an example of a null-recurrent birth-death process in natural scale.
@@
$\alpha_{n}=\beta_{n}=(n+1)^{\gamma},$
where $\gamma\leq 1$.
From (5.1), we have
$m_{0}=1,m_{n}=\dfrac{1}{(n+1)^{\gamma}}(n\geq 1).$
@@
    If $\gamma=1$, then $F(N)\sim N\log N.$ By the property of Lambert W-function, $F^{-1}(t)\sim\dfrac{t}{\log t}.$ If $w(t)=t^{a}$ where $0<a<1$, then by Corollary 5.3,
    $\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{F^{-1}(t)}dt\sim\int^{\infty}\dfrac{\log t}{t^{2-a}}dt.$
@@
  2. (2)
    If $\gamma<1$, then $F(N)\sim N^{2-\gamma}$ and $F^{-1}(t)\sim t^{\frac{1}{2-\gamma}}.$ If $w(t)=t^{a}$ where $0<a<1$, then by Corollary 5.3,
    $\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{F^{-1}(t)}dt\sim\int^{\infty}t^{a-1+\frac{1}{\gamma-2}}dt.$
    Hence $E(\tau^{a})<\infty$ if and only if $a<\dfrac{1}{2-\gamma}$.

### v2

@@ 2 Preliminaries on continuous-time Markov chains
Let $T_{y}=\inf\{t>0:X_{t}=y\}$ be the first hitting time of state $y$. Notice that, if the process $X_{t}$ starts from $y$, then $P^{y}(T_{y}=0)=1$. Let the first discontinuous point be $\eta_{1}:=\inf\{t>0:X_{t}\neq X_{0}\}$. The first returning time of state $x$ is denoted by $\tau_{x}=\inf\{t>\eta_{1}:X_{t}=x\}$. It’s obviously that $P^{x}(\tau_{y}=T_{y})=1$ for $x\neq y$. According to Theorem II.15.4 in chung1967markov, both $T_{y}$ and $\tau_{x}$ have continuous probability density functions.
The transition rates of a CTMC are often described using a density matrix denoted by $Q=(q_{xy})$. Throughout this paper, we consistently assume that the Q-matrix $Q=(q_{xy})$ is regular. For a given transition probability matrix $P(t,x,y)$, a regular Q-matrix $Q=(q_{xy})$ means
@@
where $0\leq q_{xy}<\infty$ for $x\neq y$ and $0<q_{x}:=-q_{xx}<\infty,\,\sum_{y\in I}q_{xy}=0.$ However, given a Q-matrix, the process is not necessarily unique. We have already defined $\eta_{1}$ the first discontinuous point of $\{X_{t},t\geq 0\}$. By the definition of density matrix,
$P^{x}(\eta_{1}>t)=e^{-q_{x}t},$
which means the first discontinuous point of $\{X_{t},t\geq 0\}$ exists. We denote $\eta_{2}$ the first discontinuous point after $\eta_{1}$. Then $\eta_{2}$ is the second discontinuous point of $\{X_{t},t\geq 0\}$ and $\eta_{2}$ exists by the strong Markov property, and so on. From this, we know that $\eta_{n}$ is well defined. Let
$\eta:=\lim_{n\rightarrow+\infty}\eta_{n}$
be the first flying time of $\{X_{t},t\geq 0\}$. We assume throughout the paper that the first flying time is infinite, so that the process is uniquely determined by the Q-matrix.
@@
For an irreducible recurrent CTMC $\{X_{t},t\geq 0\}$ with state space $I$, the following statements are equivalent.
@@
    The invariant measure $m(\{x\})$ satisfies $m(I)=\infty$.
@@
Condition (C): $w(t)$ is a non-negative increasing function on $0,+\infty)$ and satisfies
$\lim_{t\rightarrow\infty}w^{\prime}(t)=0,\,\,w^{\prime\prime}(t)<0\mbox{ on }[t_{0},+\infty)\mbox{ for some }t_{0}>0.$ (C)
For example, $w(t)=1\vee\log^{a}t(a>0)$, $w(t)=t^{a}(0<a<1)$, $\cdots$ Intuitively speaking, a function that satisfies Condition (C) is an increasing function whose growth rate is slower than that of a linear function. Inspired by Zhao2006null, the first returning time can be characterized more precisely by considering $E^{x}(w(\tau_{x}))$ for different weight functions $w(\cdot)$ satisfying Condition (C).
@@ 3 The first returning speed of CTMCs
Fixed $T>0$. If a non-negative function $f(t)$ is integrable on $0,+\infty)$ and satisfies that
$\forall a>0,\,\int_{a}^{a+T}f(t)dt\leq\int_{0}^{T}f(t)dt,$
then
@@
By Lemma 3.1, we can give the following propositions.
@@
    Thus, by Lemma 3.1, the proof is completed.
@@
    By Lemma 3.1, we obtain the second proposition.
@@
For two non-negative measurable functions $f(t)$ and $g(t)$ on $0,+\infty)$, we say that $f(t)\sim g(t)$ if there exist constants $C_{2}>C_{1}>0$ such that $C_{1}g(t)\leq f(t)\leq C_{2}g(t)$ for all sufficiently large $t$. Furthermore, if $f(t)\sim g(t)$, then the convergence properties of $\int^{\infty}f(t)dt$ and $\int^{\infty}g(t)dt$ are the same. For such two integrals, we also write $\int^{\infty}f(t)dt\sim\int^{\infty}g(t)dt.$
The most important result in this section will be presented. We will utilize the propositions above to give an estimate of $E^{x}(w(\tau_{x}))$.
@@
By Theorem 3.3,
$\int_{k}^{\infty}\frac{t^{a-1}}{G_{1/t}(x,x)}dt=\int_{k}^{\infty}t^{a-1}\sqrt{\dfrac{1}{t}\left(\dfrac{1}{t}+4\right)}dt\sim\int^{\infty}t^{a-1}\dfrac{2}{\sqrt{t}}dt.$
Hence, for $a>0$, we have $E^{x}(\tau_{x}^{a})<\infty$ if and only if $a<1/2$.
@@
Consider a CTMC with the state space $I=\mathbf{Z}^{2}$. Each state $(m,n)$ transitions to the four adjacent lattice points $(m,n\pm 1)$ and $(m\pm 1,n)$ at a rate of $1$. it is easy to check the process is null recurrent. The embedded chain of this process is two-dimensional simple random walk.
@@
If we take $w(t)=t^{a}(0<a<1)$, then the integral in (3.5) is always divergent. Assume that $w(t)=\log^{a}t$ where $a>0$. By Theorem 3.3, consider that
@@
$=a(1-a)\pi\int^{\infty}\dfrac{(\log t)^{a-3}}{t}dt+a\pi\int^{\infty}\dfrac{(\log t)^{a-2}}{t}dt$
$\sim\int^{\infty}\dfrac{(\log t)^{a-2}}{t}dt.$
Hence, for $a>0$, we have $E^{x}(\log^{a}\tau_{x})<\infty$ if and only if $a<1$.
@@ 4 Reversible CTMCs and electric networks
Given a network $(I,E,C)$ and two boundaries $a,b\in I$, a voltage $v(x)$ is a function $v:I\rightarrow\mathbf{R}$ that is harmonic at any $x\in I\backslash\{a,b\}$. We say $v(x)$ as the unit voltage between $a$ and $b$, that is, $v$ is the voltage, and $v(a)=1,v(b)=0.$
@@
$P_{\Delta}(x,y)=\begin{cases}\dfrac{q_{x}}{q_{x}+\lambda}\dfrac{c_{xy}}{c_{x}},&x,y\neq\Delta,\\ \dfrac{\lambda}{q_{x}+\lambda},&x\neq\Delta,y=\Delta.\end{cases}$
@@ 5 The first returning speed of birth-death process
$m_{0}=1,\,m_{n}=\dfrac{\beta_{0}\beta_{1}\beta_{2}\cdots\beta_{n-1}}{\alpha_{1}\alpha_{2}\cdots\alpha_{n-1}\alpha_{n}}\mbox{\,\,for\,\,}n\geq 1,$ (5.1)
@@
$s(0)=0,\,s(1)=\beta_{0}^{-1},\,s(n)=\beta_{0}^{-1}+\sum_{i=1}^{n-1}\dfrac{\alpha_{1}\alpha_{2}\cdots\alpha_{i}}{\beta_{0}\beta_{1}\cdots\beta_{i}}\mbox{\,\,for\,\,}n\geq 2.$ (5.2)
@@
$c_{n,n+1}=m_{n}\beta_{n}=\begin{cases}\beta_{0},&n=0,\\ \dfrac{\beta_{0}\beta_{1}\beta_{2}\cdots\beta_{n}}{\alpha_{1}\alpha_{2}\cdots\alpha_{n-1}\alpha_{n}},&n\geq 1,\end{cases}$ (5.3)
@@
Since what we need to estimate is the situation when $\lambda$ is small, we can always assume $A_{1}>B_{2}$. For the electric network corresponding to a recurrent process, since the conductance $A_{n}$ should tend to 0, we have $\lim_{n\to\infty}A_{n}<\lim_{n\to\infty}B_{n}$. By Theorem 4.8, it can be seen that the closer $A_{n-1}$ is to $B_{n}$, the better the estimation of $C_{0\leftrightarrow\Delta}(\lambda)$ will be. Using this method, we present the estimation of some important models.
@@
Assume that $\{X_{t},t\geq 0\}$ is a null-recurrent birth-death process on $\mathbf{N}$. Take $w(t)$ satisfied Condition (C). Let $F(n)=M(n)s(n)$ and
@@
  1. (1)
    If
    $\limsup_{n\rightarrow\infty}\dfrac{M(n)}{M(n-1)}<\infty,$
    then $E(w(\tau))<\infty$ if and only if
    $\int^{\infty}|w^{\prime\prime}(t)|M(F^{-1}(t))dt<\infty.$
  2. (2)
    If
    $\limsup_{n\rightarrow\infty}\dfrac{s(n)}{s(n-1)}<\infty,$
    then $E(w(\tau))<\infty$ if and only if
    $\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{s(F^{-1}(t))}dt<\infty.$
@@
Take $N=\inf\{n\geq 0:A_{n-1}\leq B_{n}\}$. Since $A_{1}>B_{2}$, $A_{n}\downarrow 0$ and $B_{n}\uparrow\infty$, we have $3\leq N<\infty$ and
@@
$\dfrac{A_{N-1}}{2}\leq[A_{N-1},B_{N}]\leq C_{0\leftrightarrow\Delta}(\lambda)\leq A_{N-1}+B_{N}\leq 2B_{N},$
@@
$\dfrac{B_{N-1}}{2}\leq\dfrac{1}{2}\max(A_{N-1},B_{N-1})\leq C_{0\leftrightarrow\Delta}(\lambda)\leq 2\min\left(B_{N},A_{N-2}\right)\leq 2B_{N}.$ (5.4)
@@
then there is a constant $k>0$ such that $\dfrac{M(n)}{M(n-1)}\leq k$ for all $n\geq 1.$ Thus, we have
$\dfrac{1}{2}\leq\dfrac{C_{0\leftrightarrow\Delta}(\lambda)}{\lambda M(N-1)}=\dfrac{C_{0\leftrightarrow\Delta}(\lambda)}{B_{N-1}}\leq 2\dfrac{B_{N}}{B_{N-1}}=2\dfrac{M(N)}{M(N-1)}\leq 2k.$
Therefore, take $\lambda=1/t$ and $N=F^{-1}(\lambda^{-1})$, by Theorem 4.4,
$E(w(\tau))<\infty\Leftrightarrow\int^{\infty}|w^{\prime\prime}(t)|M(F^{-1}(t))dt<\infty.$
The second statement in the theorem can also be deduced analogously from (5.4). ∎
Here, we present an example of a null-recurrent birth-death process.
@@
$\alpha_{n}=\beta_{n}=(n+1)^{\gamma},\,\,\gamma\leq 1.$
From (5.1) and (5.2), we have $s(n)=n$ and
$m_{0}=1,\quad m_{n}=\dfrac{1}{(n+1)^{\gamma}},\,n\geq 1.$
@@
    If $\gamma<1$, then $F(N)\sim N^{2-\gamma}$ and $F^{-1}(t)\sim t^{\frac{1}{2-\gamma}}.$ Let $w(t)=t^{a}(0<a<1)$, then by Theorem 5.1,
    $\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{s(F^{-1}(t))}dt\sim\int^{\infty}t^{a-1+\frac{1}{\gamma-2}}dt.$
    Hence $E(\tau^{a})<\infty$ if and only if $a<\dfrac{1}{2-\gamma}$.
  2. (2)
    If $\gamma=1$, then $F(N)\sim N\log N.$ By the property of Lambert W-function, $F^{-1}(t)\sim\dfrac{t}{\log t}.$ Let $w(t)=t^{a}(0<a<1)$, then by Theorem 5.1,
    $\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{s(F^{-1}(t))}dt\sim\int^{\infty}\dfrac{\log t}{t^{2-a}}dt.$
@@
    If we take $w(t)=\dfrac{t}{\log^{b}t}(b>0)$, then we have
    $\int^{\infty}\dfrac{t|w^{\prime\prime}(t)|}{s(F^{-1}(t))}dt\sim\int^{\infty}\dfrac{1}{t\log^{b}t}dt.$
    Hence for $b>0$, $E(w(\tau))<\infty$ always holds for $b>1$.